# UniBERT: Adversarial Training for Language-Universal Representations

### v2

@@ 5.5 Discussion
Figure 3: Relative mean improvements in UniBERT and other models when trained on individual languages (i.e. "each") as opposed to all languages (i.e. "all").
@@
The results indicate that the multilingual training strategy, in conjunction with the adversarial objective, has a pronounced impact on overall performance. Figure 3 illustrates the relative improvements achieved by each model when transitioning from individual language fine-tuning (denoted as "each") to joint multilingual fine-tuning (denoted as "all") [^6] Since the score on the STS is between 0 and 1, we normalized it before calculating the mean relative improvement.. mBERT-base, Distil-mBERT-base, and the XLM-RoBERTa versions exhibit a moderate average relative improvement of approximately 1.17%: 1.51% for Distil-mBERT-base, 0.62% for mBERT-base, 2.11% for XLM-RoBERTa-base, and 0.44% for XLM-RoBERTa-large. In contrast, UniBERT models indicate significantly greater gains, with an average relative improvement of 7.72%: 3.76% for UniBERT-small, 10.33% for uniBERT-XSmall, and 9.08% for UniBERT-XXSmall.
@@
This significant difference suggests that the adversarial training component, which promotes the learning of language-agnostic representations, is particularly effective when combined with joint multilingual training. For example, the UniBERT-XSmall model improves from 40.20% in the "each" configuration to 44.35% in the "all" configuration, and the UniBERT-XXSmall model similarly increases from 34.51% to 37.64%. Such improvements are notably higher than those observed in larger models, where the increase is relatively modest (e.g., XLM-RoBERTa-base and XLM-RoBERTa-large improve by less than 1.2 and 0.3 percentage points, respectively).
@@
Furthermore, a statistical t-test resulted in a t-statistic of 2.715 with a p-value of 0.0181 when comparing the relative improvements between the UniBERT models (12 entries, average improvement of 7.72%) and the other models (16 entries, average improvement of 1.17%). The observed differences are statistically significant because the p-value is lower than the significance threshold of $\alpha=0.05$. This analysis underscores that combining joint multilingual training with the adversarial objective enhances absolute performance and significantly benefits models with limited capacity by effectively leveraging cross-lingual representations.
@@ 6 Conclusion
In this paper, we provide UniBERT, a multilingual language model that employs a novel training framework consisting of masked language modeling, adversarial training, and knowledge distillation. Our extensive evaluations of four NLP tasks (i.e., named entity recognition, natural language inference, question answering, and semantic textual similarity) demonstrate that the proposed approach achieves competitive performance relative to established baselines and benefits significantly from joint multilingual training. The UniBERT variants, for example, obtained an average relative improvement of 7.72% between tasks. However, the larger baseline models showed an improvement of 1.17%. Integrating the adversarial objective has proven particularly effective in fostering language-invariant representations, enhancing cross-lingual generalization, and producing statistically significant gains (i.e., p-value = 0.0181) over models trained in individual languages when trained jointly on multilingual data.
The promising results of UniBERT underscore the potential of our methodology in addressing the challenges associated with multilingual processing in resource-constrained settings. Although our model achieves a favorable balance between efficiency and performance, there remains ample scope for future work, including further refinement of adversarial strategies, exploration of alternative distillation techniques, and extension to additional languages and domains. Overall, the contributions of this work provide a robust foundation for future research in efficient multilingual modeling and adaptable and scalable NLP systems.
7 Limitations
Despite the promising results achieved by UniBERT, our approach has certain limitations, primarily due to computational constraints that influenced our model design choices. To ensure feasibility with available resources, we opted for compact architectures with fewer parameters than the larger models used for comparison. Although this design allows efficiency and scalability, it limits the model’s capacity to learn and retain complex linguistic representations. As a result, UniBERT exhibits lower absolute performance, particularly in tasks that require deep contextual understanding, such as question answering and semantic textual similarity. However, our findings indicate that joint multilingual training and adversarial learning effectively compensate for reduced capacity, allowing UniBERT to achieve competitive results despite its smaller size.
Although adversarial training and multilingual learning help mitigate the drawbacks of a compact model, they do not eliminate the inherent limitations of reduced parameterization. Larger models retain an advantage in absolute performance due to their ability to capture more intricate linguistic structures. However, our methodology remains fundamentally sound, demonstrating that efficient multilingual training strategies can yield strong cross-lingual generalization even with fewer resources. This balance between efficiency and performance makes UniBERT a compelling solution for resource-constrained applications. Future work could explore scaling UniBERT with additional computational resources to close the performance gap further while maintaining its efficiency benefits.

### v3

@@ 5.6 Discussion
Figure 3: Relative improvements in UniBERT and other models when trained on individual languages (i.e., case "each") as opposed to all languages (i.e., case "all").
@@
The results indicate that the multilingual training strategy, in conjunction with the adversarial objective, has a pronounced impact on overall performance. Figure 3 illustrates the relative improvements achieved by each model when transitioning from individual language fine-tuning (denoted as "each") to multilingual joint fine-tuning (denoted as "all") [^6] Since the score on the STS is between 0 and 1, we normalized it before calculating the mean relative improvement.. mBERT-base, Distil-mBERT-base, and the XLM-RoBERTa versions exhibit a moderate average relative improvement of approximately 1.12%: 1.51% for Distil-mBERT-base, 0.62% for mBERT-base, 2.11% for XLM-RoBERTa-base, 0.44% for XLM-RoBERTa-large, and 0.95% for XLM-V-base. In contrast, UniBERT models indicate significantly greater gains, with an average relative improvement of 7.72%: 3.76% for UniBERT-small, 10.33% for uniBERT-XSmall, and 9.08% for UniBERT-XXSmall.
@@
This significant difference suggests that the adversarial training component, which promotes the learning of language-agnostic representations, is particularly effective when combined with multilingual joint training. For example, the UniBERT-XSmall model improves from 40.20% in the "each" configuration to 44.35% in the "all" configuration, and the UniBERT-XXSmall model similarly increases from 34.51% to 37.64%. Such improvements are notably higher than those observed in larger models, where the increase is relatively modest (e.g., XLM-RoBERTa-base and XLM-RoBERTa-large improve by less than 1.2 and 0.3 percentage points, respectively).
Performance variations between model sizes can be attributed to several key factors. Architecturally, UniBERT models with fewer parameters face capacity limitations when trained in individual languages, creating greater potential for improvement through cross-lingual knowledge sharing. Larger models like XLM-RoBERTa already effectively capture language-specific features during individual training, limiting additional gains from joint training. The adversarial objective functions differently between model sizes. In smaller architectures, it serves as an effective regularize, preventing overfitting to language-specific patterns while promoting efficient parameter usage. The Wikipedia training corpus provides consistent cross-lingual patterns that smaller models leverage more effectively with adversarial guidance. This effect is most pronounced in tasks requiring deeper semantic processing, such as NLI and QA, where UniBERT-XSmall demonstrates relative improvements of 12.7% and 28.5% respectively, suggesting adversarial training helps compact models prioritize cross-lingual semantic features over language-specific patterns.
@@
Furthermore, a statistical t-test resulted in a t-statistic of 2.714 with a p-value of 0.0184 when comparing the relative improvements between the UniBERT models (12 entries, average improvement of 7.72%) and the other models (16 entries, average improvement of 1.12%). The observed differences are statistically significant because the p-value is lower than the significance threshold of $\alpha=0.05$. This analysis underscores that combining multilingual joint training with the adversarial objective enhances absolute performance and significantly benefits models with limited capacity by effectively leveraging cross-lingual representations.
The domain adaptation capabilities of UniBERT models further validate our approach. When evaluated in the legal domain using the MAPA dataset, UniBERT models demonstrate remarkable improvements when trained jointly across languages, with relative gains of 38.16% for UniBERT-Small, 51.36% for UniBERT-XSmall, and 55.71% for UniBERT-XXSmall. This suggests that language-invariant representations learned through adversarial training facilitate effective cross-domain knowledge transfer, particularly benefiting specialized domains where annotated data may be scarce for specific languages.
6 Limitations and Future Work
Despite the promising results achieved by UniBERT, our approach faces certain limitations, primarily stemming from computational constraints that influenced our model design choices. To ensure feasibility with available resources, we opted for compact architectures with fewer parameters than the larger models used for comparison. This design choice, while enabling efficiency and scalability, inherently limits the model’s capacity to learn and retain complex linguistic representations. Consequently, UniBERT exhibits lower absolute performance, particularly in tasks requiring deep contextual understanding, such as question answering and semantic textual similarity.
Our current mitigation strategy is primarily based on multilingual joint training and adversarial learning, which have proven effective in compensating for reduced capacity. These techniques promote the learning of language-agnostic representations that generalize well across languages, allowing UniBERT to achieve competitive results despite its smaller size. However, these approaches do not eliminate the inherent limitations of reduced parameterization. Larger models maintain an advantage in absolute performance due to their ability to capture more intricate linguistic structures and relationships.
Several additional technical strategies could further address these limitations without necessarily increasing the size of the model. One promising approach is to implement more sophisticated knowledge distillation techniques, such as layer-wise distillation with attention matching sajedi2023datadam , which could enable better transfer of representational capabilities from larger teacher models. Parameter-efficient fine-tuning methods, such as adapter modules, LoRA (Low-Rank Adaptation) hulora , or BitFit zaken2022bitfit , represent viable strategies to enhance task-specific performance while maintaining a compact base model. Furthermore, implementing targeted bias mitigation strategies pozzi2025mitigating would help ensure fairness across languages without expanding the model parameters, potentially reducing representational disparities between high- and low-resource languages while maintaining the compact nature of the model.
Future technical research could pursue two complementary directions: further model compression and performance optimization. Additional model compression techniques, such as quantization shao2024dq , pruning corrHanMD15 , and specialized knowledge distillation variants for multilingual settings, could further reduce the model size while preserving cross-lingual capabilities. Specialized fine-tuning distillation protocols targeting specific linguistic phenomena across language families zhou2023universalner ; he2025study could substantially improve performance in tasks that require deeper semantic understanding. Another promising research direction would involve systematically evaluating compact multilingual models on edge devices across various hardware configurations, such as measurements of inference latency, memory usage, and power consumption in real-world deployment scenarios, particularly in multilingual mobile applications.
Beyond technical improvements, future work could investigate broader ethical and practical considerations of compact multilingual models. Researchers may develop robust bias mitigation strategies specifically designed for multilingual contexts, including methods to identify and remediate language-specific biases that may be amplified during model compression. Studies examining representation disparities between high- and low-resource languages in compressed models compared to their larger counterparts would provide valuable insights into fairness issues. Furthermore, expanding training datasets beyond Wikipedia to include more diverse text sources and genres across all 107 languages, with particular emphasis on increasing representation of low-resource languages, could address fundamental data limitations. Such an expansion would benefit from careful data curation processes to ensure quality and balance, potentially incorporating linguistic annotations to enhance the model’s understanding of language-specific phenomena.
@@ 7 Conclusion
In this paper, we introduce UniBERT, a multilingual language model that employs a novel training framework consisting of masked language modeling, adversarial training, and knowledge distillation. Our extensive evaluations of four NLP tasks (i.e., named entity recognition, natural language inference, question answering, and semantic textual similarity) demonstrate that the proposed approach achieves competitive performance relative to established baselines and benefits significantly from multilingual joint training. The UniBERT variants, for example, obtained an average relative improvement of 7.72% between tasks. However, the larger baseline models showed an improvement of 1.12%.
Integrating the adversarial objective has proven particularly effective in fostering language-invariant representations, enhancing cross-lingual generalization, and producing statistically significant gains (i.e., p-value = 0.0184) over models trained in individual languages when trained jointly on multilingual data. Furthermore, our out-of-domain evaluation of legal documents demonstrates that these language-invariant representations facilitate cross-domain generalization, with UniBERT models showing up to 55.71% relative improvement when trained jointly across languages.
The promising results of UniBERT underscore the potential of our methodology in addressing the challenges associated with multilingual processing in resource-constrained settings. Although our model balances efficiency and performance, there remains ample scope for future work, including further refinement of adversarial strategies, exploration of alternative distillation techniques, and extension to additional languages and domains. Overall, the contributions of this work provide a robust foundation for future research in efficient multilingual modeling and adaptable and scalable NLP systems.